# Scaling Robust Optimization for Multi-Agent Robotic Systems: A Distributed Perspective

### v1

@@ VIII-B Proposition 2 Proof
VIII-C Proposition 3 Proof
The proof of Proposition 3 is similar to the one of Proposition 2. First, let us rewrite constraint (47) as follows
@@ VIII-C Proposition 3 Proof
where $h_{k,\bar{m}}^{i}{}^{\mathrm{T}},h_{k,\bar{m}}^{j}{}^{\mathrm{T}}$ are the $\bar{m}^{th}$ rows of the matrices ${\bf A}_{i}{\bf P}_{k}^{i},{\bf A}_{j}{\bf P}_{k}^{j}$ respectively. By combining (VIII-C) and (VIII-C), we obtain
@@
$\|{\bm{\mu}}_{d,k}^{i}+{\bm{\mu}}_{d,k}^{j}\|_{2}\leq\tilde{c}_{ij}-c_{ij}.$ (110)
VIII-D Distributed Architecture Global Updates
@@
$\{{\bm{\nu}}^{\bar{u},l+1}_{i},\;{\bm{\nu}}^{d,l+1}_{i}\}_{i=1}^{N}$
$~{}~{}~{}=\operatornamewithlimits{argmin}\;\pazocal{L}_{\rho_{\bar{u}},\rho_{d}}(\{\tilde{{\bm{\mu}}}^{\bar{u},l+1}_{i},\tilde{{\bm{\mu}}}^{d,l+1}_{i},\tilde{{\bm{\nu}}}^{\bar{u}}_{i},\tilde{{\bm{\nu}}}^{d}_{i},{\bm{\lambda}}_{i}^{\bar{u},l},{\bm{\lambda}}_{i}^{d,l}\}_{i=1}^{N})$
$~{}~{}~{}=\operatornamewithlimits{argmin}\;\sum_{i=1}^{N}{\bm{\lambda}}_{i}^{\bar{u}}{}^{\mathrm{T}}(\tilde{{\bm{\mu}}}^{\bar{u}}_{i}-\tilde{{\bm{\nu}}}^{\bar{u}}_{i})+{\bm{\lambda}}_{i}^{d}{}^{\mathrm{T}}(\tilde{{\bm{\mu}}}^{d}_{i}-\tilde{{\bm{\nu}}}^{d}_{i})$ (111)
$~{}~{}~{}~{}\qquad\qquad+\frac{\rho_{\bar{u}}}{2}\|\tilde{{\bm{\mu}}}^{\bar{u}}_{i}-\tilde{{\bm{\nu}}}^{\bar{u}}_{i}\|_{2}^{2}+\frac{\rho_{d}}{2}\|\tilde{{\bm{\mu}}}^{d}_{i}-\tilde{{\bm{\nu}}}^{d}_{i}\|_{2}^{2}$
@@
${\bm{\nu}}^{\bar{u},l+1}_{i},{\bm{\nu}}^{d,l+1}_{i}$ (112)
$~{}~{}~{}=\operatornamewithlimits{argmin}_{{\bm{\nu}}^{\bar{u}}_{i},{\bm{\nu}}^{d}_{i}}\;\bar{{\bm{\lambda}}}_{i}^{\bar{u},l}{}^{\mathrm{T}}(\hat{{\bm{\mu}}}^{\bar{u},l+1}_{i}-{\bm{\nu}}^{\bar{u}}_{i})+\bar{{\bm{\lambda}}}_{i}^{d,l}{}^{\mathrm{T}}({\bm{\mu}}^{d,l+1}_{i}-{\bm{\nu}}^{d}_{i})$
$~{}~{}~{}\qquad\qquad+\frac{\rho_{\bar{u}}}{2}\|\hat{{\bm{\mu}}}^{\bar{u},l+1}_{i}-{\bm{\nu}}^{\bar{u}}_{i}\|_{2}^{2}+\frac{\rho_{d}}{2}\|{\bm{\mu}}^{d,l+1}_{i}-{\bm{\nu}}^{d}_{i}\|_{2}^{2}$
$~{}~{}~{}\qquad\qquad+\sum_{\hat{j}\in\pazocal{P}_{i}}\big{(}\bar{{\bm{\lambda}}}_{\hat{j}}^{i,\bar{u},l}{}^{\mathrm{T}}(\bar{{\bm{\mu}}}^{i,\bar{u},l+1}_{\hat{j}}-{\bm{\nu}}^{\bar{u}}_{i})$
$~{}~{}~{}\qquad\qquad+\bar{{\bm{\lambda}}}_{\hat{j}}^{i,d,l}{}^{\mathrm{T}}(\bar{{\bm{\mu}}}^{i,d,l+1}_{\hat{j}}-{\bm{\nu}}^{d}_{i})+\frac{\rho_{\bar{u}}}{2}\|\bar{{\bm{\mu}}}^{i,\bar{u},l+1}_{\hat{j}}-{\bm{\nu}}^{\bar{u}}_{i}\|_{2}^{2}$
$~{}~{}~{}\qquad\qquad+\frac{\rho_{d}}{2}\|\bar{{\bm{\mu}}}^{i,d,l+1}_{\hat{j}}-{\bm{\nu}}^{d}_{i}\|_{2}^{2}\big{)}$
To find the minimum on the RHS, let us differentiate the RHS with respect to the global variables ${\bm{\nu}}^{\bar{u}}_{i},{\bm{\nu}}^{d}_{i}$ and equate it to zero. For ${\bm{\nu}}^{\bar{u},l+1}_{i}$, we get
$-\bar{{\bm{\lambda}}}_{i}^{\bar{u},l}+\rho_{\bar{u}}(\hat{{\bm{\mu}}}^{\bar{u},l+1}_{i}-{\bm{\nu}}^{\bar{u}}_{i})$ (113)
$\qquad+\sum_{\hat{j}\in\pazocal{P}_{i}}\big{(}-\bar{{\bm{\lambda}}}_{\hat{j}}^{i,\bar{u},l}+\rho_{\bar{u}}(\bar{{\bm{\mu}}}^{i,\bar{u},l+1}_{\hat{j}}-{\bm{\nu}}^{\bar{u}}_{i})\big{)}=0$
@@
${\bm{\nu}}^{\bar{u},l+1}_{i}$ $=\frac{1}{(n(\pazocal{P}_{i})+1)}\bigg{(}\frac{\bar{{\bm{\lambda}}}_{i}^{\bar{u},l}}{\rho_{\bar{u}}}+\hat{{\bm{\mu}}}_{i}^{\bar{u},l+1}$ (114)
$\qquad\qquad+\sum_{\hat{j}\in\pazocal{P}_{i}}\big{(}\frac{\bar{{\bm{\lambda}}}_{\hat{j}}^{i,\bar{u},l}}{\rho_{\bar{u}}}+\bar{{\bm{\mu}}}_{\hat{j}}^{i,\bar{u},l+1}\big{)}\bigg{)}$
The update for ${\bm{\nu}}^{d,l+1}_{i}$ can be derived similarly.
IX Simulation Data
IX-A System Parameters

### v2

@@ B-B Proposition 2 Proof
B-C Proposition 3 Proof
The proof of Proposition 3 is similar to the one of Proposition 2. First, let us rewrite constraint ( 40 ) as follows
@@ B-C Proposition 3 Proof
where $h_{k,\bar{m}}^{i}{}^{\mathrm{T}},h_{k,\bar{m}}^{j}{}^{\mathrm{T}}$ are the $\bar{m}^{th}$ rows of the matrices ${\bf A}_{i}{\bf P}_{k}^{i},{\bf A}_{j}{\bf P}_{k}^{j}$ respectively. By combining (98) and (100), we obtain
@@
$\|{\bm{\mu}}_{d,k}^{i}+{\bm{\mu}}_{d,k}^{j}\|_{2}\leq\tilde{c}_{ij}-c_{ij}$ (105)
B-D Interpretation of tighter approximation in 2D space
Here, we provide the intuition behind the approximation in a 2D case considering that the non-convex norm constraints correspond to the obstacle avoidance constraints. In that case, ${\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}$ represents the disturbance-free position of the agent $i$, and $\bm{b}_{i}$ represents the center of the obstacle.
We considered the following two approximations -
B-D1 First Approximation
Using the triangle inequality, we considered the following approximation. The obstacle avoidance constraint
$\min_{\bm{\zeta}^{i}\in\pazocal{U}_{i}}\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}+{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}\bm{\zeta}^{i}-\bm{b}_{i}\|_{2}\geq c_{i}$ (106)
is approximated by the following constraint
$\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}-\bm{b}_{i}\|_{2}-\max_{\bm{\zeta}^{i}\in\pazocal{U}_{i}}\|{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}\bm{\zeta}^{i}\|_{2}\geq c_{i}.$ (107)
Let us now analyze each of the above constraints. Since, the disturbance $\bm{\zeta}^{i}$ lies in an ellipse, and the mean is a linear transformation of $\bm{\zeta}^{i}$, the possible realizations of the state mean at each time step would also form an ellipse. Fig.1 and Fig.2 illustrate a two-agent case with an obstacle (black circle) and the mean state of two agents $i$ and $j$ represented by the ellipses. In Fig. 11, $d_{min}^{i}$ and $d_{min}^{j}$ represent the smallest distance between the obstacle and the state mean realizations of the agents $i$ and $j$, respectively. The RHS of the original constraint (106), i.e., $\min_{\bm{\zeta}^{i}\in\pazocal{U}_{i}}\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}+{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}\bm{\zeta}^{i}-\bm{b}_{i}\|_{2}$, represents $d_{min}^{i}$. Thus, the original constraint is to have $d_{min}^{i}\geq c_{i}$.
Now, the RHS of the tighter approximate constraint (107) involves two terms: $\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}-\bm{b}_{i}\|_{2}$ and $\max_{\bm{\zeta}^{i}\in\pazocal{U}_{i}}\|{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}\bm{\zeta}^{i}\|_{2}$. The term $\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}-\bm{b}_{i}\|_{2}$ represents the distance between the disturbance-free state of the agent $i$ (represented by the corresponding center of ellipse in Fig 11) and the obstacle, and the term $\max_{\bm{\zeta}^{i}\in\pazocal{U}_{i}}\|{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}\bm{\zeta}^{i}\|_{2}$ is the maximum possible deviation (represented by a black arrow in Fig 11b) in the state mean due to the deterministic disturbance. Thus, the tighter approximation is analogous to considering that the deviation in the state mean is the same in all directions, with magnitude equal to $\max_{\bm{\zeta}^{i}\in\pazocal{U}_{i}}\|{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}\bm{\zeta}^{i}\|_{2}$. As shown in Fig. 11b, $\hat{d}_{min}^{i}$ represents the smallest distance between the boundary of the circle formed with the center as disturbance-free state’s mean ${\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}$ and radius equal to the maximum deviation $\max_{\bm{\zeta}^{i}\in\pazocal{U}_{i}}\|{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}\bm{\zeta}^{i}\|_{2}$. Thus, the original constraint $d_{min}^{i}\geq c_{i}$ is approximated to $\hat{d}_{min}^{i}\geq c_{i}$. Further, it can be observed that the approximation only affects the agent $i$ and not the agent $j$.
Figure 11: Demonstration of first approximation of obstacle avoidance constraint: Robust mean trajectory realizations of two agents represented as ellipses.
B-D2 Second Approximation
The second type of approximation is based on the fact that
$\max_{{\bm{y}}\in\mathscr{Y}}\|{\bf A}{\bm{y}}\|_{2}$ $=\max_{{\bm{y}}\in\mathscr{Y}}\bigg{(}\sum_{\bar{m}=1}^{m}({\bm{a}}_{\bar{m}}^{\mathrm{T}}{\bm{y}})^{2}\bigg{)}^{1/2}$ (108)
$\leq\bigg{(}\sum_{\bar{m}=1}^{m}\max_{{\bm{y}}\in\mathscr{Y}}({\bm{a}}_{\bar{m}}^{\mathrm{T}}{\bm{y}})^{2}\bigg{)}^{1/2}$ (109)
for some matrix ${\bf A}\in\mathbb{R}^{m\times n_{y}}$, and vector ${\bm{y}}\in\mathbb{R}^{n_{y}}$. The constraint
is approximated by
$\|{\bm{\mu}}_{d,k}^{i}{}\|_{2}\leq\tilde{c}_{i}-c_{i}.$ (111)
with the introduction of the slack variable ${\bm{\mu}}_{d,k}^{i}$. It should be noted that the value of ${\bm{\mu}}_{d,k}^{i}$ should provide a good approximation to the maximum state deviation term $\max_{\bm{\zeta}^{i}\in\pazocal{U}_{i}}\|{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}\bm{\zeta}^{i}\|_{2}$ when the agent $i$ is closer to the obstacle. Since ${\bm{\mu}}_{d,k}^{i}$ is a slack variable, $(\tau^{i})^{1/2}\|\bm{\Gamma}_{i}^{\mathrm{T}}{\bf M}_{i}^{\mathrm{T}}h_{k,\bar{m}}^{i}\|_{{\bf S}_{i}^{-1}}=({\bm{\mu}}_{d,k}^{i})_{\bar{m}}$ during these instants. Fig. 12 provides various examples of realizations of the state mean. The black arrows in the ellipses in Fig. 12a represent the maximum deviation in the state mean. The red arrows in Fig 12b represent the quantity $\|{\bm{\mu}}_{d,k}^{i}\|_{2}$ with each element $({\bm{\mu}}_{d,k}^{i})_{\bar{m}}=(\tau^{i})^{1/2}\|\bm{\Gamma}_{i}^{\mathrm{T}}{\bf M}_{i}^{\mathrm{T}}h_{k,\bar{m}}^{i}\|_{{\bf S}_{i}^{-1}}$. It can be observed that $\|{\bm{\mu}}_{d,k}^{i}\|_{2}$ provides a good approximation of the maximum deviation for different shapes of the state mean realizations.
Figure 12: Demonstration of second type of approximation: Effect of the approximation on various robust mean trajectory realization shapes.
B-E Distributed Architecture Global Updates
@@
$\{\tilde{{\bm{\nu}}}_{\bar{u}}^{i,l+1},\tilde{{\bm{\nu}}}_{d}^{i,l+1}\}_{i\in\pazocal{V}}$
$~{}~{}~{}=\operatornamewithlimits{argmin}\;\pazocal{L}_{\rho_{\bar{u}},\rho_{d}}(\{\tilde{{\bm{\mu}}}_{\bar{u}}^{i,l+1},\tilde{{\bm{\mu}}}_{d}^{i,l+1},\tilde{{\bm{\nu}}}_{\bar{u}}^{i},{\tilde{{\bm{\nu}}}_{d}^{i},{\bm{\lambda}}^{i,l}_{\bar{u}},{\bm{\lambda}}^{i,l}_{d}}\}_{i\in\pazocal{V}})$
$~{}~{}~{}=\operatornamewithlimits{argmin}\;\sum_{i=1}^{N}{\bm{\lambda}}^{i,l}_{\bar{u}}{}^{\mathrm{T}}(\tilde{{\bm{\mu}}}_{\bar{u}}^{i,l+1}-\tilde{{\bm{\nu}}}_{\bar{u}}^{i})+{\bm{\lambda}}^{i,l}_{d}{}^{\mathrm{T}}(\tilde{{\bm{\mu}}}_{d}^{i,l+1}-\tilde{{\bm{\nu}}}_{d}^{i})$ (112)
$~{}~{}~{}~{}\qquad\qquad+\frac{\rho_{\bar{u}}}{2}\|\tilde{{\bm{\mu}}}_{\bar{u}}^{i,l+1}-\tilde{{\bm{\nu}}}_{\bar{u}}^{i}\|_{2}^{2}+\frac{\rho_{d}}{2}\|\tilde{{\bm{\mu}}}_{d}^{i,l+1}-\tilde{{\bm{\nu}}}_{d}^{i}\|_{2}^{2}$
@@
${\bm{\nu}}_{\bar{u}}^{i,l+1},{\bm{\nu}}_{d}^{i,l+1}$ (113)
$~{}~{}~{}=\operatornamewithlimits{argmin}_{{\bm{\nu}}_{\bar{u}}^{i},{\bm{\nu}}_{d}^{i}}\;\bar{{\bm{\lambda}}}^{i,l}_{\bar{u}}{}^{\mathrm{T}}(\hat{{\bm{\mu}}}_{\bar{u}}^{i,l+1}-{\bm{\nu}}_{\bar{u}}^{i})+\bar{{\bm{\lambda}}}^{i,l}_{d}{}^{\mathrm{T}}({\bm{\mu}}_{d}^{i,l+1}-{\bm{\nu}}_{d}^{i})$
$~{}~{}~{}\qquad\qquad+\frac{\rho_{\bar{u}}}{2}\|\hat{{\bm{\mu}}}_{\bar{u}}^{i,l+1}-{\bm{\nu}}_{\bar{u}}^{i}\|_{2}^{2}+\frac{\rho_{d}}{2}\|{\bm{\mu}}_{d}^{i,l+1}-{\bm{\nu}}_{d}^{i}\|_{2}^{2}$
$~{}~{}~{}\qquad\qquad+\sum_{\hat{j}\in\pazocal{P}_{i}}\big{(}\bar{{\bm{\lambda}}}^{\hat{j},l}_{i,\bar{u}}{}^{\mathrm{T}}(\bar{{\bm{\mu}}}_{i,\bar{u}}^{\hat{j},l+1}-{\bm{\nu}}_{\bar{u}}^{i})$
$~{}~{}~{}\qquad\qquad+\bar{{\bm{\lambda}}}^{\hat{j},l}_{i,d}{}^{\mathrm{T}}(\bar{{\bm{\mu}}}_{i,d}^{\hat{j},l+1}-{\bm{\nu}}_{d}^{i})+\frac{\rho_{\bar{u}}}{2}\|\bar{{\bm{\mu}}}_{i,\bar{u}}^{\hat{j},l+1}-{\bm{\nu}}_{\bar{u}}^{i}\|_{2}^{2}$
$~{}~{}~{}\qquad\qquad+\frac{\rho_{d}}{2}\|\bar{{\bm{\mu}}}_{i,d}^{\hat{j},l+1}-{\bm{\nu}}_{d}^{i}\|_{2}^{2}\big{)}$
To find the minimum on the RHS, let us differentiate the RHS with respect to the global variables ${\bm{\nu}}_{\bar{u}}^{i},{\bm{\nu}}_{d}^{i}$ and equate it to zero. For ${\bm{\nu}}_{\bar{u}}^{i,l+1}$, we get
$-\bar{{\bm{\lambda}}}^{i,l}_{\bar{u}}+\rho_{\bar{u}}(\hat{{\bm{\mu}}}_{\bar{u}}^{i,l+1}-{\bm{\nu}}_{\bar{u}}^{i})$ (114)
$\qquad+\sum_{\hat{j}\in\pazocal{P}_{i}}\big{(}-\bar{{\bm{\lambda}}}^{\hat{j},l}_{i,\bar{u}}+\rho_{\bar{u}}(\bar{{\bm{\mu}}}_{i,\bar{u}}^{\hat{j},l+1}-{\bm{\nu}}_{\bar{u}}^{i})\big{)}=0$
@@
${\bm{\nu}}_{\bar{u}}^{i,l+1}$ $=\frac{1}{(n(\pazocal{P}_{i})+1)}\bigg{(}\frac{\bar{{\bm{\lambda}}}^{i,l}_{\bar{u}}}{\rho_{\bar{u}}}+\hat{{\bm{\mu}}}^{i,l+1}_{\bar{u}}$ (115)
$\quad\quad\quad+\sum_{\hat{j}\in\pazocal{P}_{i}}\big{(}\frac{\bar{{\bm{\lambda}}}^{\hat{j},l}_{i,\bar{u}}}{\rho_{\bar{u}}}+\bar{{\bm{\mu}}}^{\hat{j},l+1}_{i,\bar{u}}\big{)}\bigg{)}$
The update for ${\bm{\nu}}_{d}^{i,l+1}$ can be derived similarly.
B-F Proposition 4 Proof
We derive the complexity of solving each local subproblem of an agent in the distributed SDP framework versus the proposed distributed framework. The computational burden of the SDP and SOCP constraints is significant compared to the linear constraints. Therefore, we consider these constraints to determine the worst-case complexity of the algorithms. Further, the complexity of solving an optimization problem with SDP or SOCP constraints [25] depends on the number of variables ($n_{\text{var}}$), the number of SDP or SOCP constraints ($n_{\text{SDP}}$ or $n_{\text{SOCP}}$), and the size of each SDP or SOCP constraint involved in the optimization problem. Let us also assume that $n_{c}$ represents the size of $c^{th}$ constraint.
B-F1 Complexity of the Distributed SDP Framework
For this framework. we assume that the shared variables among the agents are the control parameters since it is not straightforward to have state means as the shared variables in the distributed SDP approach. Now, let us write down the total number of variables involved in the sub-problem of agent $i$ as follows
$n_{\text{var}}$ $=T\big{(}n_{u_{i}}+n_{u_{i}}\gamma_{h}n_{x_{i}}+n_{obs}+n_{\text{inter}}$ (116)
$~{}~{}~{}~{}\quad\quad\quad+\sum_{j\in\pazocal{N}_{i}}n_{u_{j}}+n_{u_{j}}\gamma_{h}n_{x_{j}}\big{)}$
where $T(n_{u_{i}}+n_{u_{i}}\gamma_{h}n_{x_{i}})$, $T(\sum_{j\in\pazocal{N}_{i}}n_{u_{j}}+n_{u_{j}}\gamma_{h}n_{x_{j}})$ correspond to the control parameters of agent $i$ and its neighbor agents $j\in\pazocal{N}_{i}$ respectively, and $T(n_{obs}+n_{\text{inter}})$ correspond to the dual variables (as in $\beta$) in the collision avoidance constraints in (28). Further, this framework involves SDP constraints whose details are given as follows
  • For obstacle avoidance: number of constraints = $Tn_{obs}$, size of each constraint = $\bar{n}_{i}+1$.
  • For inter-agent collision avoidance: number of constraints = $Tn_{\text{inter}}$, size of each constraint = $\bar{n}_{i}+\bar{n}_{j}+1$.
The complexity of each iteration of an interior-point method to solve the local sub-problem with the SDP constraints is given as $O(n_{\text{var}}^{2}\sum_{c=1}^{n_{\text{SDP}}}n_{c}^{2})$ [25]. First, let us rewrite $\sum_{c=1}^{n_{\text{SDP}}}n_{c}^{2}$ as follows -
$\sum_{c=1}^{n_{\text{SDP}}}n_{c}^{2}$ $=Tn_{\text{obs}}(\bar{n}_{i}+1)^{2}+Tn_{\text{inter}}(\bar{n}_{i}+\bar{n}_{j}+1)^{2}$ (117)
As we are interested in finding $O(n_{\text{var}}^{2}\sum_{c=1}^{n_{\text{SDP}}}n_{c}^{2})$, let us simplify the contribution of $n_{\text{var}}^{2}$ and $\sum_{c=1}^{n_{\text{SDP}}}n_{c}^{2}$ to eliminate the lesser order terms in $n_{\text{var}}^{2}\sum_{c=1}^{n_{\text{SDP}}}n_{c}^{2}$. Assuming all the agents have the same control and state dimensions, and are equal to $n_{u_{i}}$ and $n_{x_{i}}$ respectively, $n_{\text{var}}^{2}$ can be simplified to
$T^{2}\big{(}(n_{\text{inter}}+1)n_{u_{i}}\gamma_{h}n_{x_{i}}+n_{\text{obs}}\big{)}^{2}$ (118)
Next, we assume that $\bm{\Gamma}_{i}=I$, which gives us $\bar{n}_{i}=Tn_{d_{i}}+n_{x_{i}}$. Using this, $\sum_{c=1}^{n_{\text{\text{SDP}}}}n_{c}^{2}$ yields
$T\big{(}n_{\text{obs}}T^{2}n_{d_{i}}^{2}+n_{\text{inter}}(T^{2}n_{d_{i}}^{2}+T^{2}n_{d_{j}}^{2}+T^{2}n_{d_{i}}n_{d_{j}})\big{)}$ (119)
Assuming $n_{d_{i}}=n_{d_{j}}$, we could further simplify the above to the following.
$(n_{\text{obs}}+n_{\text{inter}})T^{3}n_{d_{i}}^{2}$ (120)
By combining the above simplifications, we get the complexity per interior point iteration to be
$O\Big{(}T^{2}\big{(}(n_{\text{inter}}+1)n_{u_{i}}$ $\gamma_{h}n_{x_{i}}+n_{obs}\big{)}^{2}$ (121)
$~{}~{}~{}~{}\quad(n_{obs}+n_{\text{inter}})T^{3}n_{d_{i}}^{2}\Big{)}$
which can be upper bounded by
$O\Big{(}T^{5}(n_{\text{inter}}+n_{obs})^{3}n_{u_{i}}^{2}n_{x_{i}}^{2}n_{d_{i}}^{2}\gamma_{h}^{2}\Big{)}$ (122)
Let $L_{\text{IP}}$ be the number of interior point iterations taken to reach the desired convergence. Then, the total complexity of solving the sub-problem is given as
$O\Big{(}L_{\text{IP}}T^{5}(n_{\text{inter}}+n_{obs})^{3}n_{u_{i}}^{2}n_{x_{i}}^{2}n_{d_{i}}^{2}\gamma_{h}^{2}\Big{)}.$ (123)
B-F2 Complexity of the Proposed Distributed Framework
Similar to the previous derivation, let us first write the expression for $n_{\text{var}}$ involved in the sub-problem of agent $i$. We have
$n_{\text{var}}$ $=T(n_{u}+n_{u}\gamma_{h}n_{x}+m(1+2n_{\text{inter}})$ (124)
$\qquad\qquad\qquad+n_{\text{obs}}+n_{\text{inter}})$
where $T(n_{u}+n_{u}\gamma_{h}n_{x})$ corresponds to the control parameters of agent $i$, $Tm(1+2n_{\text{inter}})$ corresponds to copy variables of neighbor agents at agent $i$, and $T(n_{obs}+n_{\text{inter}})$ corresponds to the slack variables $\tilde{c}_{i}$, $\tilde{c}_{ij}$ in the collision avoidance constraints. This framework involves SOCP constraints whose details are given as follows
  • For obstacle avoidance (constraint (34)): number of constraints = $Tn_{obs}$, size of each constraint = $m+1$.
  • For inter-agent collision avoidance (constraint (39)): number of constraints = $Tn_{\text{inter}}$, size of each constraint = $m+1$.
  • Bounds on ${\bm{\mu}}_{d}^{i}$ (constraint (35)): number of constraints = $Tm$, size of each constraint = $\bar{n}_{i}+1$.
Firstly, it should be noted that the constraints of type (35) are of significantly larger size than the constraints of type (34) and (39). The constraints (35) depend neither on the number of obstacles nor on the number of neighbor agents. Second, all the above constraints are SOCP constraints and are computationally less expensive than solving the SDP constraints of the same size. The complexity of each iteration of an interior-point method to solve the local sub-problem with the SOCP constraints is given as $O(n_{\text{var}}^{2}\sum_{c=1}^{n_{\text{SOCP}}}n_{c})$ [25]. Let us now rewrite $\sum_{c=1}^{n_{\text{SOCP}}}n_{c}$ as follows
$\sum_{c=1}^{n_{\text{SOCP}}}n_{c}$ $=Tn_{\text{obs}}(m+1)+Tn_{\text{inter}}(m+1)$ (125)
$\qquad\qquad\qquad\qquad\qquad+Tm(\bar{n}_{i}+1).$
Based on the same assumptions as in the previous derivation, we simplify the contribution of each term in $O(n_{\text{var}}^{2}\sum_{c=1}^{n_{\text{SOCP}}}n_{c})$ starting with $n_{\text{var}}^{2}$ simplified to the following
$T^{2}(n_{u}\gamma_{h}n_{x}+mn_{\text{inter}}+n_{\text{obs}})^{2}$ (126)
since $m$ is constant. We can further reduce the above to
$T^{2}(n_{u}\gamma_{h}n_{x}+n_{\text{inter}}+n_{\text{obs}})^{2}.$ (127)
Now, the contribution of $\sum_{c=1}^{n_{\text{SOCP}}}n_{c}$ can be written as
$T(mn_{\text{obs}}+mn_{\text{inter}}+mTn_{d_{i}})$ (128)
which can be further simplified to
$T(n_{\text{obs}}+n_{\text{inter}}+Tn_{d_{i}}).$ (129)
Combining the above simplifications for $n_{\text{var}}^{2}$ and $\sum_{c=1}^{n_{\text{SOCP}}}n_{c}$, we have the computational complexity per interior point iteration for solving the local subproblem given as
$O\Big{(}T^{3}(n_{u}\gamma_{h}n_{x}+n_{\text{inter}}+n_{\text{obs}})^{2}\big{(}n_{\text{obs}}+n_{\text{inter}}+Tn_{d_{i}}\big{)}\Big{)}$ (130)
which could be upper bounded by
$O\Big{(}T^{4}(n_{\text{inter}}+n_{obs})^{3}n_{u_{i}}^{2}n_{x_{i}}^{2}n_{d_{i}}\gamma_{h}^{2}\Big{)}$ (131)
If $L_{\text{IP}}$ is the number of interior point iterations taken to reach the optimal value, the total complexity of solving the sub-problem is given by
$O\Big{(}L_{\text{IP}}T^{4}(n_{\text{inter}}+n_{obs})^{3}n_{u_{i}}^{2}n_{x_{i}}^{2}n_{d_{i}}\gamma_{h}^{2}\Big{)}.$ (132)
Appendix C Simulation Data
C-A System Parameters